The speed of light is 3 x 10 8 m/s. Convert this to furlongs and fortnight. A furlong is 220 yards, and a fortnight is 14 days. An inch is 2.54 cm


1
Expert's answer
2021-08-31T08:58:26-0400
"1m = 1.09361\space yards = \dfrac{1.09361}{220}\space furlongs"

"1s = \dfrac{1}{60\cdot 60\cdot 24} days = \dfrac{1}{60\cdot 60\cdot 24\cdot 14} fortnight"

Thus, obtain:


"c = 3\times 10^8\dfrac{m}{s} = 3\times 10^8\cdot \dfrac{\dfrac{1.09361}{220}\space furlongs}{\dfrac{1}{60\cdot 60\cdot 24\cdot 14} fortnight} =\\
=3\times 10^8\cdot \dfrac{1.09361}{220}\cdot 60\cdot 60\cdot 24\cdot 14\dfrac{ furlongs}{fortnight} =\\
\approx 1.8\times 10^{12}\dfrac{ furlongs}{fortnight}"

Answer. "1.8\times 10^{12}\dfrac{ furlongs}{fortnight}".